# Memories of Ian G. Macdonald

### v1

@@ Abstract
This is a slightly edited translation of a paper in Dutch which appeared in Nieuw Archief voor Wiskunde (5) 25 (2024), No.2, 87–90 on the occasion of I. G. Macdonald’s death in 2023, and aimed at a very broad mathematical audience. First we review some of Macdonald’s most important older results. Then we focus on the period 1985–1995 when Macdonald often visited the Netherlands and there was much interaction between his work, notably the Macdonald polynomials, and the work by the authors. We end with some glimpses about Macdonald as a person.
@@
Ian Grant Macdonald was an English mathematician, known for his contributions in algebra, geometry, Lie theory, combinatorics and special functions. He passed away on August 8, 2023 at the age of 94. See obituaries by Peter Cameron [2] and Arun Ram [35]. Ram also wrote a longer memorial paper [36]. Much earlier A. O. Morris wrote a nice biographical paper [29] about him. In the present memorial article we will in particular focus on our contacts with Macdonald starting in the eighties of the previous century.
Macdonald studied at Trinity College in Cambridge, where he graduated in 1952. Instead of subsequently preparing a PhD thesis, he worked during five years as a civil servant. But then he could no longer resist to return to university, where he had consecutive appointments in Manchester, Exeter, Oxford and London. He was elected Fellow of the Royal Society in 1979. In 2002 he received an honorary doctorate of the University of Amsterdam. In his word of thanks he expressed his great satisfaction to finally having received the doctor’s degree.
Below we will first discuss some highlights of Macdonald’s work in the seventies of the previous century. Next we turn to his work starting in the middle of the eighties on orthogonal polynomials associated with root systems. It was in particular this topic where Macdonald’s work had a close relationship with ours. There was a mutual stimulation, by letters and by discussions during the regular visits which he paid to our country. What follows below is therefore not only a survey of Macdonald’s work but also a piece of history of Dutch mathematics.
Photo credits: First two photos by Peter van Emde Boas, last two photos by Margrit Rösler and Michael Voit.
@@
For a better understanding of Macdonald’s work one needs to have some notion of a root system. Let $V$ be a Euclidean vector space of dimension $n$ where vectors $\alpha,\beta$ have inner product $(\alpha,\beta)$. For a nonzero vector $\alpha\in V$ we denote by
@@
The theory of root systems and their symmetry groups is a chapter in Euclidean geometry that can be well understood by a second year undergraduate student in mathematics with some knowledge of linear algebra and group theory. The irreducible root systems, i.e., those which cannot be written as an orthogonal union of two smaller ones, are classified by the symbols
@@
and accompanying Coxeter diagrams. By the work of Claude Chevalley in the fifties of the last century we know that they are also the language of the combinatorial infrastructure of the simple algebraic groups $\mathbf{G}(\mathbb{F})$ over a field $\mathbb{F}$. The representation theory of these groups leads to subtle identities in terms of root systems. Conversely such identities make the representation theory computable.
@@
The first result is the $p$-adic analogue of the elementary spherical functions on $\mathbf{G}(\mathbb{Q}_{p})/\mathbf{G}(\mathbb{Z}_{p})$, announced by Macdonald [21] in 1968. The details appeared in a book [22] which is based on lectures at the Ramanujan Institute in Madras in 1970 and which is written in a very accessible way. One could get a free copy by just sending a request to the author. In a sense, the $p$-adic case is more simple than Harish-Chandra’s real case, because the elementary spherical functions for the $p$-adics are given by an explicit formula (for $\mathbf{G}=\mathbf{GL}_{n}$ in terms of Hall–Littlewood polynomials [26, §V.3]). Macdonald’s explicit formula was one of the ingredients used by Robert Langlands for the formulation of the conjectures named after him [17, p.39].
For a good understanding of the second result we first say somewhat more about Weyl’s character formula. This gives the value of the character on a maximal torus as a quotient of two Fourier polynomials, both skew-symmetric under the Weyl group $W$. The expression in the numerator contains a parameter $\lambda$ (Élie Cartan’s highest weight) which characterizes the irreducible representation. The denominator is the famous Weyl denominator $\Delta$, explicitly given by
$\Delta=\prod_{\alpha>0}(e^{\alpha/2}-e^{-\alpha/2}),$
where the product runs over the set of positive roots $\alpha>0$, i.e., the roots lying on one side of a fixed generic linear subspace of $V$ of codimension 1. When we apply the character formula to the trivial representation, the quotient expresses the trivial character $1$. Hence the numerator specialised at $\lambda=0$ equals the Weyl denominator $\Delta$, which is the celebrated Weyl denominator formula.
In this formula of the form $\prod$=$\sum$ = ∑ the product has terms with opposite signs which cancel each other, by which less terms remain in the sum than first thought. Macdonald observed a similar phenomenon in Jacobi’s triple product identity
@@
This is a classical formula in terms of theta functions in which $q,z$ are complex variables (with $z\neq 0$ and $|q|<1$). In this formula Macdonald recognized the structure of the affine root system of type $\widetilde{\mathrm{A}}_{1}$. Next he found analogous identities for all other affine root systems [23]. By specialization he also obtained simple summation formulas for $\eta^{d}$, where Dedekind’s eta function is given by
@@
Macdonald felt a special sympathy for the general linear group $\mathbf{GL}_{n}$ and the symmetric group $S_{n}$, the “All-embracing Majesties” of the linear algebraic and the finite groups, respectively (as Weyl [40, end of Ch. IV] phrased it for $\mathbf{GL}_{n}$ in ironic adoration). Here the representation theory has a strongly combinatorial flavour, with methods which are, in spite of their elementary nature, not necessarily simple. In 1979 he published the influential book Symmetric Functions and Hall Polynomials on this matter. A much expanded second edition [26] appeared in 1995. For this work he received in 2009 the Leroy P. Steele Prize for Mathematical Exposition.
@@
Almost parallel, in 1987, Macdonald had circulated handwritten manuscripts which described the polynomials now known as Macdonald polynomials. His first manuscript treated a class of symmetric polynomials with two parameters $q$ and $t$. These extrapolated both the Jack polynomials and the Hall–Littlewood polynomials by adding a parameter $q$. He also gave an explicit system of commuting difference operators having his polynomials as eigenfunctions. In the same year Simon Ruijsenaars [37] had independently found this explicit system. Macdonald gives an extensive treatment of this class of symmetric polynomials in [26, Ch. VI]. This chapter also includes the so-called Pieri formula [26, §VI.6] for these polynomials, which Koornwinder [15] proved in 1988 by showing a duality between the variables and the spectral parameters. For this duality Koornwinder was inspired by the two-variable case, where the (homogeneous) Macdonald polynomials are essentially one-variable continuous $q$-ultraspherical polynomials. There he could immediately observe the duality, familiar as he was with the $q$-polynomial stuff by the work of his former teacher Dick Askey [1].
Macdonald’s next two manuscripts in 1987 concerned $q$-analogues of Jacobi polynomials associated with root systems, now known as Macdonald polynomials associated with root systems. They are the direct $q$-analogues of the polynomials introduced by Heckman and Opdam in the same time, which now could be seen as $q\to 1$ limits of Macdonald’s polynomials. For root system of type A Macdonald’s $q$-polynomials are contained in his class of symmetric functions, so it was after the first manuscript a natural step to pass to general root systems. Possibly the work of Heckman en Opdam, which he had closely witnessed by several visits to Leiden in 1987 and 1988, has also been a motivation for him to look at $q$-analogues of their polynomials. The $q$-parameter in Macdonald’s $q$-polynomials gave an additional freedom by which he could prove full orthogonality in an elementary way. And thus, taking the limit for $q\to 1$, the full orthogonality of the Heckman–Opdam polynomials now also had an elementary proof. Macdonald also gave conjectures for the norm and the evaluation at unity of his polynomials. In the limit case $q\to 1$ his conjectures were in accordance with the results of Heckman and Opdam. In 1988 Macdonald put together the two manuscripts into one manuscript [25]. There he also mentioned Opdam’s thesis and proof of the constant term conjecture.
@@
In the fall of 1989 the Berlin wall was turned down, and in the summer of the next year many Russians visited the ICM in Kyoto, among whom there was Ivan Cherednik. In a few years all conjectures of Macdonald were made theorems by Cherednik [3]. It was not always easy reading in Ivan’s papers, but expositions such as by Opdam [34] and Macdonald [28] were helpful for understanding. The key to all this were Hecke algebras: first the affine Hecke algebra of Iwahori and Matsumoto coming from the $p$-adic world, next the degenerate Hecke algebra of Drinfeld [5], and for the duality between variable and spectral parameter the double affine Hecke algebra of Cherednik. The commuting generalized Dunkl operators had non-symmetric (non-$W$-invariant) Macdonald polynomials as eigenfunctions, and these were projected onto the symmetric ($W$-invariant) Macdonald polynomials by the symmetrization map.
Another new development were the Koornwinder polynomials [16]. These are a significant 5-parameter extension of the 3-parameter family of Macdonald polynomials associated with the non-reduced root system of type $\mathrm{BC}_{n}$ (here $q$ is not considered as a parameter). This was at the same time an $n$-variable extension of the $4$-parameter family of Askey–Wilson-polynomials [1] in one variable. Also for this case Macdonald formulated conjectures, which again could be proved by Hecke algebra methods [38], [28]. Younger Dutch mathematicians were stimulated by these developments, see van Diejen [4] and Stokman [39] (former students of Ruijsenaars and Koornwinder, respectively). Earlier van der Lek wrote an influential thesis [18]. This was helpful for Ion [12] to give a conceptual proof of the duality theorem for double affine Hecke algebras which was proved in Macdonald’s last book [28, Theorem (3.5.1)] using a case-by-case analysis.
@@
We conclude with some words about Ian Macdonald as a person and as a mathematician. He was a very gifted speaker. During his lecture in Séminaire Bourbaki in the middle of the nineties [27] one of us was sitting beside Jim Arthur, an expert in the Langlands program. At the end of the lecture he was silent for a while, and then said: “What an outstanding expositor”. We felt indeed privileged that Macdonald visited the Netherlands so frequently to share his views with us. On these occasions Ian was usually accompanied by his wife Greta, originating from the Flemish part of Belgium. Ian had no driving license, and thus she drove while he read the map for navigation. During a lecture in the early nineties Ian started as follows:
@@
No doubt Ian was a mathematician with a very classical inclination. During his education in Cambridge the spirit of Hardy and Ramanujan may still have been felt. Ian was a master in dealing with formulas, but in retrospect all these formulas were directed to a better understanding of the representation theory of the simple algebraic groups (over a field or ring). We are truly grateful that the mathematical road travelled by Ian in his later life crossed so frequently with ours.

### v2

@@ Abstract
This is a slightly edited translation of a paper in Dutch which appeared in Nieuw Archief voor Wiskunde (5) 25 (2024), No.2, 87–90 on the occasion of I. G. Macdonald’s death in 2023, and aimed at a very broad mathematical audience. First we review some of Macdonald’s most important older results. Then we focus on the period 1985–1995, when Macdonald often visited the Netherlands and there was much interaction between his work, notably the Macdonald polynomials, and the work by the authors. We end with some glimpses about Macdonald as a person.
@@
Ian Grant Macdonald was an English mathematician, known for his contributions in algebra, geometry, Lie theory, combinatorics and special functions. He passed away on August 8, 2023 at the age of 94. See obituaries by Peter Cameron [3] and Arun Ram [36]. Ram also wrote a longer memorial paper [37]. Much earlier A. O. Morris has written a nice biographical paper [30] about him. In the present memorial article we will in particular focus on our contacts with Macdonald starting in the 1980s.
Macdonald studied at Trinity College in Cambridge, where he graduated in 1952. Instead of subsequently preparing a PhD thesis, he worked for five years as a civil servant. But then he could no longer resist the urge to return to academia, where he had consecutive appointments in Manchester, Exeter, Oxford and London. He was elected Fellow of the Royal Society in 1979. In 2002 he received an honorary doctorate from the University of Amsterdam. In his word of thanks he expressed his great satisfaction to finally having received the doctor’s degree.
Below we will first discuss some highlights of Macdonald’s work in the 1970s. Next we turn to his work starting in the mid 1980s on orthogonal polynomials associated with root systems. It was in particular this topic where Macdonald’s work had a close relationship with ours. There was a mutual stimulation, by letters and by discussions during the regular visits which he paid to our country. What follows below is therefore not only a survey of Macdonald’s work but also a piece of history of Dutch mathematics.
@@
For a better understanding of Macdonald’s work one needs to have some notion of a root system. The theory of root systems and their symmetry groups [2] is a chapter in Euclidean geometry that can be well understood by a second year undergraduate student in mathematics with some knowledge of linear algebra and group theory. We give the basic definitions.
Let $V$ be a Euclidean vector space of dimension $n$ where vectors $\alpha,\beta$ have inner product $(\alpha,\beta)$. For a nonzero vector $\alpha\in V$ we denote by
@@
The irreducible root systems, i.e., those which cannot be written as an orthogonal union of two smaller ones, are classified by the symbols
@@
and accompanying Coxeter diagrams. By the work of Claude Chevalley in the 1950s, we know that they are also the language of the combinatorial infrastructure of the simple algebraic groups $\mathbf{G}(\mathbb{F})$ over a field $\mathbb{F}$. The representation theory of these groups leads to subtle identities in terms of root systems. Conversely such identities make the representation theory computable.
@@
The first result is the $p$-adic analogue of the elementary spherical functions on $\mathbf{G}(\mathbb{Q}_{p})/\mathbf{G}(\mathbb{Z}_{p})$, announced by Macdonald [22] in 1968. The details appeared in a book [23] which is based on lectures at the Ramanujan Institute in Madras in 1970 and which is written in a very accessible way. In the old days, one could get a free copy by just sending a request to the author. In a sense, the $p$-adic case is simpler than Harish-Chandra’s real case, because the elementary spherical functions for the $p$-adics are given by an explicit formula (for $\mathbf{G}=\mathbf{GL}_{n}$ in terms of Hall–Littlewood polynomials [27, §V.3]). Macdonald’s explicit formula was one of the ingredients used by Robert Langlands for the formulation of the conjectures named after him [18, p.39].
For a good understanding of the second result [24] we first say somewhat more about Weyl’s character formula. This gives the value of the character on a maximal torus as a certain quotient. The denominator is a Fourier polynomial, skew-symmetric under the Weyl group $W$ and containing a parameter $\lambda$ (Élie Cartan’s highest weight) which characterizes the irreducible representation. The denominator is the famous Weyl denominator $\Delta=\prod_{\alpha>0}(e^{\alpha/2}-e^{-\alpha/2})$, where the product runs over the set of positive roots $\alpha>0$, i.e., the roots lying on one side of a fixed generic linear subspace of $V$ of codimension 1. When we apply the character formula to the trivial representation, the quotient expresses the trivial character $1$. Hence the numerator specialised at $\lambda=0$ equals the Weyl denominator $\Delta$, which is the celebrated Weyl denominator formula
$\prod_{\alpha>0}(e^{\alpha/2}-e^{-\alpha/2})=\sum_{w\in W}\epsilon(w)e^{w(\rho)},$
where $\epsilon(w)=\det(w)$ and $\rho$ is half the sum of the positive roots. In this formula of the form $\prod$=$\sum$ = ∑ the product has terms with opposite signs which cancel each other, so that fewer terms remain in the sum than first thought.
Macdonald observed a similar phenomenon in Jacobi’s triple product identity
@@
This is a classical formula in terms of theta functions in which $q,z$ are complex variables (with $z\neq 0$ and $|q|<1$). In this formula Macdonald recognized the structure of the affine root system of type $\mathrm{A}_{1}$. Next he found analogous identities for all other affine root systems [24]. By specialization he also obtained simple summation formulas for $\eta^{d}$, where Dedekind’s eta function is given by
@@
Macdonald felt a special sympathy for the general linear group $\mathbf{GL}_{n}$ and the symmetric group $S_{n}$, the “All-embracing Majesties” of the linear algebraic and the finite groups, respectively (as Weyl [41, end of Ch. IV] phrased it for $\mathbf{GL}_{n}$ in ironic adoration). Here the representation theory has a strongly combinatorial flavour, with methods which are, in spite of their elementary nature, not necessarily simple. In 1979 he published the influential book Symmetric Functions and Hall Polynomials on this matter. A much expanded second edition [27] appeared in 1995. For this work he received in 2009 the Leroy P. Steele Prize for Mathematical Exposition.
@@
Almost parallel, in 1987, Macdonald had circulated handwritten manuscripts which described the polynomials now known as Macdonald polynomials. His first manuscript treated a class of symmetric polynomials with two parameters $q$ and $t$. These extrapolated both the Jack polynomials and the Hall–Littlewood polynomials by adding a parameter $q$. He also gave an explicit system of commuting difference operators having his polynomials as eigenfunctions. In the same year Simon Ruijsenaars [38] had independently found this explicit system. Macdonald gives an extensive treatment of this class of symmetric polynomials in [27, Ch. VI]. This chapter also includes the so-called Pieri formula [27, §VI.6] for these polynomials, which Koornwinder [16] proved in 1988 by showing a duality between the variables and the spectral parameters. For this duality Koornwinder was inspired by the two-variable case, where the (homogeneous) Macdonald polynomials are essentially one-variable continuous $q$-ultraspherical polynomials. There he could immediately observe the duality, familiar as he was with the $q$-polynomial stuff by the work of his former teacher Dick Askey [1].
Macdonald’s next two manuscripts in 1987 concerned polynomials now known as Macdonald polynomials associated with root systems. They are the direct $q$-analogues of Jacobi polynomials associated with root systems as introduced by Heckman and Opdam in the same time. Their polynomials could now be seen as $q\to 1$ limits of Macdonald’s polynomials. For the root system of type A, Macdonald’s $q$-polynomials are contained in his class of symmetric functions, so it was after the first manuscript a natural step to pass to general root systems. Possibly the work of Heckman and Opdam, which he had closely witnessed by several visits to Leiden in 1987 and 1988, has also been a motivation for him to look at $q$-analogues of their polynomials. The $q$-parameter in Macdonald’s $q$-polynomials gave an additional freedom by which he could prove full orthogonality in an elementary way. And thus, taking the limit for $q\to 1$, the full orthogonality of the Heckman–Opdam polynomials now also had an elementary proof. Macdonald also gave conjectures for the norm and the evaluation at unity of his polynomials. In the limit case $q\to 1$ his conjectures were in accordance with the results of Heckman and Opdam. In 1988 Macdonald put together the two manuscripts into one manuscript [26]. There he also mentioned Opdam’s thesis and proof of the constant term conjecture.
@@
In the fall of 1989 the Berlin wall was torn down, and in the summer of the next year many Russians visited the ICM in Kyoto, among whom there was Ivan Cherednik. In a few years all conjectures of Macdonald were made theorems by Cherednik [4]. It was not always easy reading in Ivan’s papers, but expositions such as by Opdam [35] and Macdonald [29] were helpful for understanding. The key to all this were Hecke algebras: first the affine Hecke algebra of Iwahori and Matsumoto coming from the $p$-adic world, next the degenerate Hecke algebra of Drinfeld [6], and for the duality between variable and spectral parameter the double affine Hecke algebra of Cherednik. The commuting generalized Dunkl operators had non-symmetric (non-$W$-invariant) Macdonald polynomials as eigenfunctions, and these were projected onto the symmetric ($W$-invariant) Macdonald polynomials by the symmetrization map.
Another new development were the Koornwinder polynomials [17]. These are a significant 5-parameter extension of the 3-parameter family of Macdonald polynomials associated with the non-reduced root system of type $\mathrm{BC}_{n}$ (here $q$ is not considered as a parameter). This was at the same time an $n$-variable extension of the $4$-parameter family of Askey–Wilson-polynomials [1] in one variable. Also for this case Macdonald formulated conjectures, which again could be proved by Hecke algebra methods [39], [29]. Younger Dutch mathematicians were stimulated by these developments, see van Diejen [5] and Stokman [40] (former students of Ruijsenaars and Koornwinder, respectively). Earlier van der Lek wrote an influential thesis [19]. This was helpful for Ion [13] to give a conceptual proof of the duality theorem for double affine Hecke algebras which was proved in Macdonald’s last book [29, Theorem (3.5.1)] using a case-by-case analysis.
@@
We conclude with some words about Ian Macdonald as a person and as a mathematician. He was a very gifted speaker. During his lecture in Séminaire Bourbaki in the mid 1990s [28] one of us was sitting beside Jim Arthur, an expert in the Langlands program. At the end of the lecture he was silent for a while, and then said: “What an outstanding expositor”. We felt indeed privileged that Macdonald visited the Netherlands so frequently to share his views with us. On these occasions Ian was usually accompanied by his wife Greta, originating from the Flemish part of Belgium. Ian had no driving license, and thus she drove while he read the map for navigation. During a lecture in the early 1990s Ian started as follows:
@@
No doubt Ian was a mathematician with a very classical inclination. During his education in Cambridge, the spirit of Hardy and Ramanujan may still have been felt. Ian was a master in dealing with formulas, but in retrospect all these formulas were directed to a better understanding of the representation theory of the simple algebraic groups (over a field or ring). We are truly grateful that the mathematical road travelled by Ian in his later life crossed so frequently with ours.
Acknowledgements
The first two photos are by Peter van Emde Boas, and the last two ones by Margrit Rösler and Michael Voit. We thank the referee for suggesting various textual improvements.